OpenStudy (anonymous):

Can someone please check this What is the value of x? Round your answer to the nearest tenth. x^2 = 86 i got ±86 but i don't know if that's correct

OpenStudy (anonymous):

What is the value of y? 4y^2 + 8 = 264 i got ± 8

OpenStudy (solomonzelman):

no (for the question in blue)

OpenStudy (solomonzelman):

when you have \(\large\color{black}{ x^2=86 }\) to get the answer you take the square root of both sides, not just of x. \(\large\color{black}{ x^2=86 }\) \(\large\color{red}{\sqrt{ \color{black}{x^2}}\color{black}{=}\pm\sqrt{\color{black}{86} } }\) see

OpenStudy (solomonzelman):

so the square root of x^2 is just x, and square root of 86 is ?

OpenStudy (solomonzelman):

(use a calculator for this, if you want a decimal value)

OpenStudy (anonymous):

9.2736184955

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

so it's 9.2?

OpenStudy (anonymous):

@iGreen